OpenStudy (anonymous):

Find the slope of line that passing through the pair of points (17,9)(5,29)

OpenStudy (anonymous):

-5/3

OpenStudy (nenadmatematika):

use formula (y2-y1)/(x2-x1)....you can call the first point A(17,9) where A(x1,y1) and the second point is some B(5,29) where B(x2,y2)...so you'll get (29-9)/(5-17)=20/(-12)=-5/3...do you understand this?